How do I write the equation of the line w/ slope 2/3 that passes through (4,-5)? I need to give the answer in slope intercept form.

Found 2 solutions by Alan3354, ewatrrr:
Answer by Alan3354(69443)

Use y = mx + b and the given point to find b. b is the y-intercept. You know the slope, it's 2/3

Answer by ewatrrr(24785)

 
Hi,


line w/ slope 2/3 that passes through (4,-5)
y = (2/3)x + b |Using ordered pair (4,-5) to solve for b
-5 = (2/3)*4 + b
-15/3 = 8/3 + b
-23/3 = b
y = (2/3)x - 23/3